OpenStudy (anonymous):

The thickness of photoresist applied to wafers in semiconductor manufacturing at a particular on the wafer is uniformly distributed between 0.2035 and 0.2161 micrometers. a) Determine the proportion of wafers that exceeds 0.2125 micrometers of photoresist thickness

OpenStudy (perl):

Did you try to use a uniform distribution for this problem?

OpenStudy (anonymous):

i tried this, \[f(x)=\frac{ 1 }{ (0.2161-0.2035 )}\] \[f(x)=\frac{ 5000 }{ 63 }\] but i don't know the limits for the integral and i'm not sure if this is correct

OpenStudy (perl):

the limit is from .2125 to the end of the domain .2161

OpenStudy (anonymous):

\[\int\limits_{0.2125}^{0.2161}\frac{ 5000 }{ 63 }\] the answer is 0.2857

OpenStudy (perl):

that's right

OpenStudy (anonymous):

Yes! Thank you so much! :)